# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 3. Gluing of Cauchy data and horizon exclusion
(34) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=-\frac{2|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{2}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}-\frac{C}{a^{2}}$
since $\mbox{tr}\chi(u_{\infty},0)\approx\mbox{tr}\chi(u_{\infty},\epsilon)$ up to $O(|u_{\infty}|^{-2})$ error terms by the choice of initial data at $u=u_{\infty}$. This provides us with the exact conditions for the non-existence of an apparent horizon in $\widehat{\mathcal{M}}_{int}$ and a dynamical existence of an apparent horizon in $J^{+}(\widehat{\mathcal{M}}_{int})\cap M_{t=-a+\epsilon}$. These are equivalent to the following condition with $\epsilon=O(a^{-\frac{1}{2}})$
(35) $\frac{3\pi(\text{Rad}(\widehat{\mathcal{M}}_{int})-\epsilon^{\frac{1}{2}})}{[\text{Rad}(\widehat{\mathcal{M}}_{int})]^{2}}+\frac{C}{a^{2}}<-\frac{2|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{2}{a}\int_{u_{\infty}}^{a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}<\frac{3\pi}{\text{Rad}(\widehat{\mathcal{M}}_{int})}+\frac{C}{a^{2}}$
Therefore, if the initial outgoing shear and the magnitude of the incoming expansion at the past null inifnity $u=u_{\infty}$ is sufficiently large, then the boundary $\partial\bigg(J^{+}(\widehat{\mathcal{M}}_{int})\cap\mathcal{M}_{t=-a+\epsilon}\bigg)=S_{-a,\epsilon}$ would have sufficiently large generalized mean curvature $H-|\kappa|$ and therefore according to the condition of [38], an apparent horizon will exist in the interior of $J^{+}(\widehat{\mathcal{M}}_{int})\cap\mathcal{M}_{t=-a+\epsilon}$.
Note the scalings of $\underline{\hat{\chi}}$ and $\mbox{tr}\underline{\chi}$ (adapted from [5])
(36) $|\underline{\hat{\chi}}(u,0)|\lesssim a^{\frac{1}{2}}|u|^{-2},~|\mbox{tr}\underline{\chi}(u_{\infty},0)|\lesssim|u_{\infty}|^{-1}$
which seem to be well-suited for the propagation of this scaling norm hierarchy. The remaining task is then to prove the semi-global existence result in the double null framework. In addition, with this scaling, note that the term $\frac{1}{a}\int_{u_{\infty}}^{-a}||u^{{}^{\prime}}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime}}$ and the non-linear contributions in the $\nabla_{3}$ equation for $\mbox{tr}\underline{\chi}$ (equation 12) contributes at the same order as the non-linear error terms ($O(a^{-2})$). Therefore, the initial large-scale-invariant norm of $\underline{\hat{\chi}}$ should be prescribed such that it dominates that of the remaining Ricci coefficients (as in 39). The solution to the subsequent Cauchy problem for t=at=-a to $t=-a+\epsilon$ for sufficiently small $\epsilon$ is a routine process and, as such, incurs significantly less difficulty. Of course, one main task is to actually construct the asymptotically flat Cauchy data on the entire slice $\mathcal{M}_{t=-a}$ from the characteristic initial development and the interior radius condition. The two main points that we want to recall are the following initial data norm
(37) $\begin{split}\Gamma^{0}_{2}:=&\sup_{S_{u,\underline{u}}\subset H_{u_{\infty}},\underline{H}_{0}}\bigg(\sum_{i\leq N}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\mbox{tr}\chi,\hat{\chi},\eta,\underline{\eta},\omega,\underline{\omega})}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\\ &+\sum_{i\leq N}\frac{a}{\lvert u\rvert^{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\mbox{tr}\underline{\chi}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\sum_{i\leq N}\frac{a}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\widetilde{\mbox{tr}\underline{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\bigg),\end{split}$
(38) $\Gamma^{0}_{\underline{\hat{\chi}}2}:=\sup_{S_{u,\underline{u}}\subset H_{u_{\infty}},\underline{H}_{0}}\sum_{i\leq N}\frac{a^{\frac{1}{2}}}{\lvert u\rvert}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\underline{\hat{\chi}}}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})},$
(39) $\Gamma^{0}_{\underline{\hat{\chi}}2}>100\Gamma^{0}_{2}$
(40) $\mathcal{R}^{0}_{2}:=\sup_{S_{u,\underline{u}}\subset H_{u_{\infty}},\underline{H}_{0}}\bigg(\sum_{i\leq N-1}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}+\sum_{i\leq N-1}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\underline{\alpha},\beta,\underline{\beta},a^{\frac{1}{2}}\rho,a^{\frac{1}{2}}\sigma)}\rVert_{L^{2}_{(sc)}(S_{u,\underline{u}})}\bigg),$
(41) $\mathcal{R}(u_{\infty},\underline{u}):=\sum_{i\leq N}\lVert{(a^{\frac{1}{2}}\nabla)^{i}\alpha}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})}+a^{\frac{1}{2}}\sum_{i\leq N}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\beta,\underline{\beta},\rho,\sigma)}\rVert_{L^{2}_{(sc)}(H_{u_{\infty}}^{(0,\underline{u})})},$
(42) $\underline{\mathcal{R}}(u,0):=\sum_{i\leq N}\bigg(\lVert{(a^{\frac{1}{2}}\nabla)^{i}\beta}\rVert_{L^{2}_{(sc)}(\underline{H}_{0}^{(u_{\infty},u)})}+a^{\frac{1}{2}}\lVert{(a^{\frac{1}{2}}\nabla)^{i}(\underline{\beta},\rho,\sigma,\underline{\alpha})}\rVert_{L^{2}_{(sc)}(\underline{H}_{\underline{u}}^{(u_{\infty},u)})}\bigg).$
Notice that this data choice is very close to the one developed by [3] and later used by [4, 11, 12] only difference being the shear. In the present framework, we need a large integrated conjugate shear $\underline{\hat{\chi}}$ on the incoming null hypersurface $\underline{H}_{0}$ while we are free to prescribe data on the initial outgoing null hypersurface $H_{u_{\infty}}$ that is compatible with the scaling 37-42. In particular, our data does not have Minkowski initial data on $\underline{H}_{0}$. In this sense, it is different from the previous choices of data considered. On the other hand, the prescription of Cauchy data on the interior part of the Cauchy slice $\mathcal{M}_{-a}$ is rather implicit and relies on the Riemannian geometric arguments of [38]. In the sequel [28], we will address the construction of the Cauchy data explicitly.
@@
Note a slight scaling difference between [3] and the current study. The $a^{\frac{1}{2}}$ gain in the decay of $\hat{\chi}$ is compensated in $\rho$.
Theorem 3.1 (Main Theorem).

### v4

@@ 2.1. Yau [44] criterion for MOTS formation
(42) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)$ $=-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime}}$
$+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},0)du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}$
and the error terms $\mathscr{E}$ decay at least $O(a^{-\frac{5}{2}})$ and therefore negligible compared to $O(a^{-1})$ term $\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}$ and $O(a^{-2})$ terms involving $\underline{\hat{\chi}}$. The two main leading order terms that we are concerned with are the $O(a^{-1})$ term and $O(a^{-2})$ term. In addition note that these estimates are stable up to at least $O(a^{-\frac{5}{2}})$ upon perturbing $u$ up to a factor $1/a$.
2.3. Interior Cauchy development and compatibility with the characteristic region
A central component of the construction is the controlled Cauchy evolution of the interior region $J^{+}(\mathcal{M}_{1})$ over a fixed time interval of length $O(1)$, starting from initial data that are pointwise small but supported on a spatial domain of diameter $\sim a$. The data are arranged so as to be compatible, along a common interface sphere, with the characteristic development constructed in Section 3, while the far exterior is completed by a Kerr end with prescribed ADM parameters. We now formulate this precisely.
Let $\mathcal{M}_{-a}$ be a smooth spacelike Cauchy hypersurface, decomposed as
$\mathcal{M}_{-a}=\mathcal{M}_{1}\;\cup\;\mathcal{M}_{2}\;\cup\;\mathcal{M}_{\mathrm{ext}},\qquad\partial\mathcal{M}_{1}=S_{-a,0},$
where:
  • $\mathcal{M}_{1}$ is a connected interior region with $H$-radius or Schoen-Yau radius comparable to $a$;
  • $\mathcal{M}_{2}\subset D_{a,1}$ is the portion contained in the characteristic development region constructed earlier, foliated by a double–null optical pair $(u,\underline{u})$ satisfying $t=u+\underline{u}=-a$ on $\mathcal{M}_{-a}$;
  • $\mathcal{M}_{\mathrm{ext}}$ denotes the asymptotic exterior region.
Fix a truncated characteristic subdomain $D^{\prime}\subset D_{a,1}$ bounded by
$u=-a-\tfrac{1}{a},\qquad\underline{u}=0,\qquad\underline{u}=1,\qquad u=u_{\infty},$
and define the enlarged interior domain
$\mathcal{M}^{1/a}_{\mathrm{int}}:=\mathcal{M}_{1}\;\cup\;\bigl(\mathcal{M}_{2}\setminus(D^{\prime}\cap\mathcal{M}_{2})\bigr).$
We prescribe smooth vacuum Cauchy data $(g,k)$ on $\mathcal{M}^{1/a}_{\mathrm{int}}$ with the following quantitative bounds: for some fixed integer $s\leq 3$,
(43) $\|k\|_{L^{\infty}(\mathcal{M}^{1/a}_{\mathrm{int}})}+\|\partial(g-\delta)\|_{L^{\infty}(\mathcal{M}^{1/a}_{\mathrm{int}})}$ $\leq C\,a^{-3/2},$
(44) $\|k\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{\mathrm{int}})}+\|\partial(g-\delta)\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{\mathrm{int}})}$ $\leq C\,a^{-3/2},$
where $H^{m}_{\mathrm{ul}}$ denotes the uniformly local Sobolev norm defined by
$\|F\|_{H^{m}_{\mathrm{ul}}(\Omega)}:=\sup_{p\in\Omega}\|F\|_{H^{m}(B_{1}(p))},$
with $B_{1}(p)$ the unit geodesic ball in the induced metric. In addition, the lapse and shift associated with the ADM decomposition satisfy
(45) $\|N-1\|_{L^{\infty}}+\|\nabla N\|_{L^{\infty}}+\|X\|_{L^{\infty}}\leq C\,a^{-3/2}\qquad\text{on }\mathcal{M}^{1/a}_{\mathrm{int}}.$
The spacetime metric in a neighborhood of $\mathcal{M}_{-a}$ is expressed in spacetime harmonic gauge, $\Box_{g}x^{\mu}=0$, and in ADM form
$\mathsfbf g=-N^{2}dt^{2}+g_{ij}\,(dx^{i}+X^{i}dt)(dx^{j}+X^{j}dt),$
so that the vacuum Einstein equations reduce to a quasilinear hyperbolic system for $g_{\mu\nu}$. In this gauge, local existence and continuation criteria depend only on uniformly local Sobolev norms of the initial data and are stable under perturbations of size $O(a^{-3/2})$ (cf. [27]).
We impose a compatibility condition across the interface sphere $S_{-a,0}=\partial\mathcal{M}_{1}$ with the data induced from the characteristic development $D_{a,1}$. Precisely, letting $(g_{\mathrm{char}},k_{\mathrm{char}})$ denote the Cauchy data induced on $\mathcal{M}_{-a}\cap D_{a,1}$, we require:
  1. (1)
    The first and second fundamental forms induced on $S_{-a,0}$ by $(\mathcal{M}_{1},g,k)$ and by $(\mathcal{M}_{2},g_{\mathrm{char}},k_{\mathrm{char}})$ agree exactly.
  2. (2)
    All tangential covariant derivatives along $S_{-a,0}$ up to order $N$ (for some fixed $N\gg 1$) coincide.
To be completely rigorous, this is accomplished by further decomposing $\mathcal{M}_{1}$ into a compact subset $\widetilde{\mathcal{M}}_{1}\subset\mathcal{M}_{1}$ and the collar region $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$. In the compact set $\widetilde{\mathcal{M}}_{1}$, we provide the data $(g,k)$ which has $H$-radius or Schoen-Yau radius $\text{Rad}_{g}$ while the data in the collar $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ is the transition region that smoothly matches the data with the characteristic development ($D_{a,1}$) induced data on $\mathcal{M}_{2}$. In particular, the collar $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ has thickness $O(\widetilde{\epsilon}a),~\widetilde{\epsilon}\in[\frac{1}{1000},\frac{1}{100}]$ and $(g,k)$ on the compact subset $\widetilde{\mathcal{M}}_{1}$ has the estimate
(46) $\|\partial^{k}k\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\leq C\,a^{-k-3/2},~\|\partial^{k+1}g\|_{L^{\infty}(\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-1}$
$\|\partial^{k}k\|_{H^{s-1}_{\mathrm{ul}}(\widetilde{\mathcal{M}}_{1})}\leq C\,a^{-k-3/2},\|\partial^{k+1}g\|_{H^{s-1}_{\mathrm{ul}}(\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-1},~k\geq 0$
where on the collar $\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1}$ the estimate (same estimate for $K$) reads
(47) $\|\partial^{k}k\|_{L^{\infty}(\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-3/2},$
(48) $\|\partial^{k}k\|_{H^{s-1}_{\mathrm{ul}}(\mathcal{M}_{1}\setminus\widetilde{\mathcal{M}}_{1})}\leq Ca^{-k-3/2},~k\geq 0$
with $\partial g$ smoothly interpolated in the collar between $\widetilde{\mathcal{M}}_{1}$ (verifying (46)) and the estimate on $\mathcal{M}_{2}$ induced by the characteric development (where in harmonic coordinates, $\partial g||\lesssim a^{-\frac{3}{2}}$) along with the same estimates for the lapse and the shift. If these are understood, if there is no further confusion, we will not separately work with this further decomposition.
@@ 2.3. Interior Cauchy development and compatibility with the characteristic region
Note that the largeness of the data is contained in the constants $C$ (which are independent of $a$ and are $O(1)$) and the smallness is encoded in the inverse power of $a$. For $a\gg 1$, the slice t=at=-a can be considered to be at a reverse late time, where one utilizes the decay ‘in time’ property of wave equations to build a large data theory where the non-linear terms are typically subdominant compared to the linear counterparts due to larger decay. This is the essential large-small control in the long time dynamics of Einstein equations.
Remark 2.
 [^4] Xuantao Chen brought this to our attention
Notice that this data verify the scale-critical estimate $\partial^{3/2}(\mathsfbf{g}-m)||_{L^{2}(\mathcal{M}_{1})}=O(1)$, $m$ being the Minkowski metric and $\mathsfbf{g}$ is the physical spacetime metric in harmonic gauge, since one derivative here costs $a^{-1}$ at best over the large domain of size $O(a)$ under study. Therefore the scale-critical norm $\dot{H}^{3/2}$ is $O(1)$ in our study. An interesting question would be to understand how small the $\dot{H}^{3/2}$ norm of $\mathsfbf{g}-m$ can be in order for the MOTS formation result to persist. This would help one understand if Minkowski space is stable under scale-critical small perturbations.
[…]
$\mathcal{M}_{1}\cup\mathcal{M}_{2}.$
Finally, we complete the data in the far exterior region $\mathcal{M}_{\mathrm{ext}}$ by a Corvino–Schoen type gluing to a Kerr initial data set outside a compact set. The glued end has ADM parameters satisfying
$m_{\mathrm{ADM}}\sim a^{1/2},\qquad|J|=O(a),$
[…]
The two analytic tasks are therefore separated as follows: (i) a controlled quasilinear hyperbolic evolution of the interior data over a time interval of length $O(1)$ (independent of $a$) in spacetime harmonic gauge, with deformation tensor and ADM variables remaining uniformly small, and (ii) a quantitative lower bound for the HH–radius along the evolved interior slices, treated in the next section.
2.4. Radius Comparison
In the previous section we showed that along the outgoing null hypersurface $H_{-a}$ the generalized mean curvature quantity $H-|\mbox{tr}_{\Sigma}K|$ is preserved up to errors of size $O(|u_{\infty}|^{-2}+a^{-5/2})$, which are negligible in the large–$a$ regime under consideration. Although monotonic decay of this quantity along outgoing null directions might be expected a priori, the admissible open class of characteristic initial data constructed here ensures that it remains effectively constant along $H_{-a}$.
The subsequent objective is to obtain quantitative control of the Schoen–Yau (or HH–) radius of the interior region of the Cauchy slice. This constitutes the central geometric step in the argument. More precisely, one must prove that the Schoen–Yau radius of the evolved interior domain
$J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}$
strictly exceeds the initial Schoen–Yau radius of $\mathcal{M}^{1/a}_{\mathrm{int}}$, and moreover to derive a sharp lower bound for this increase in terms of the large parameter $a$. We outline the geometric mechanism at a heuristic level here; the complete quantitative argument is given later, see in particular Proposition 7.2.
We refer to Figure 2. Consider first the compact manifold with boundary $\mathcal{M}_{1}$, whose boundary is the interface sphere $S_{-a,0}$. One must prescribe Cauchy data $(g,k)$ on $\mathcal{M}_{1}$ that are compatible across $S_{-a,0}$ with the data induced on the adjacent region $\mathcal{M}_{2}$ by the characteristic development in the slab $D_{a,1}$. The resulting data are then matched in the exterior region $\mathcal{M}_{3}$ with a Kerr end. This matching is achieved by a localized Corvino–Schoen type gluing construction, carried out in Section 6, which produces a global Cauchy data set on the slice $\mathcal{M}_{-a}$ satisfying the constraint equations and agreeing with the characteristic data and the Kerr data in their respective domains.
The glued initial data are then evolved from $\mathcal{M}_{-a}$ to the nearby slice $\mathcal{M}_{-a-1/a+\epsilon}$. For $\epsilon>0$ sufficiently small, depending only on the constructed initial norms, local well–posedness yields a unique spacetime development in this time interval. The data are pointwise small in the precise sense required for compatibility with the characteristic solution: the second fundamental form, shift vector field, and lapse gradient satisfy
$|k|+|X|+|\nabla N|=O(a^{-3/2})$
on $\mathcal{M}_{-a}$, while the associated ADM mass is of order $a^{1/2}$ and therefore large in the regime $a\gg 1$. In view of the quasi-linear wave equations, the allowable time of evolution may in fact be chosen uniformly of order one for example, in spacetime harmonic gauge.
Let $\Phi_{t}$ denote the flow generated by the time vector field $\partial_{t}$, which is globally defined by global hyperbolicity. Our objective is to obtain quantitative control of the transported Riemannian metric $(\Phi_{t}^{-1})^{*}g$ on the interior future domain
$J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon},$
which will be used to estimate the geometric radius and related interior quantities on the evolved slice.
During the Cauchy evolution, it is not a priori excluded that the transported metric $(\Phi_{t}^{-1})^{*}g$ on the future interior domain
$J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}$
may decrease relative to its initial size, even though $(\Phi_{t}^{-1})^{*}g$ remains quasi–isometric to the initial metric on $\mathcal{M}^{1/a}_{\mathrm{int}}$ under the smallness assumptions on the deformation tensor. At the same time, the evolved image $\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$ is a proper subset of $J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}$. Consequently, the collar region
$\Big(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}\Big)\setminus\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$
has strictly positive thickness and therefore contributes a positive amount to the geometric size measured by the Schoen–Yau (HH–) radius. The objective of this section is to show that this collar contribution dominates any possible metric contraction in the interior region; see Figure 4 for a schematic representation.
The argument relies on precise pointwise control of the second fundamental form $k$ on $\mathcal{M}_{1}$, dictated by the matching requirements with the data on $\mathcal{M}_{2}$ arising from the characteristic development in $D_{a,1}$. In the adapted frame $(T,S,e_{A})$ one has
$k_{ST}=0,\qquad|k_{SA}|+|k_{AT}|+|k_{AB}|+|k_{SS}|\;\lesssim\;a^{-3/2},$
and hence
$|k|\lesssim a^{-3/2}.$
The critical component is the tangential block $k_{AB}$, which admits the null decomposition
[…]
Individually, the traces $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ are of size $O(a^{-1})$ near the interface sphere $S_{-a,0}$ and therefore represent potentially dangerous terms. A naive estimate based solely on these trace bounds would yield $|k|=O(a^{-1})$ on $\mathcal{M}_{-a}$. Standard short–time existence theory for quasilinear wave systems would then permit a metric variation of the same order, comparable to the geometric gain coming from the collar region. Therefore, we need to turn to refined estimates.
A priori, quasilinear wave energy estimates together with direct integration of the metric transport equation along timelike and null directions suggests the comparison bound
(49) $g\Big(1-\frac{C\epsilon}{a}\Big)\;\leq\;(\Phi_{t}^{-1})^{*}g\;\leq\;g\Big(1+\frac{C\epsilon}{a}\Big)$
on the transported interior region $\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$, for a universal numerical constant $C>0$. Correspondingly, the geometric thickness contributed by the collar region
$\Big(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}\Big)\setminus\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$
is expected to satisfy a lower bound of the form
(50) $\text{Rad}_{\mathrm{collar}}\;\geq\;\text{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})\Big(1+\frac{C\epsilon}{a}\Big),$
possibly with a different constant $C$. At this level of precision, the potential metric contraction in (49) and the collar gain in (50) occur at the same order, and no effective lower bound for the radius of $J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}$ can be deduced.
The decisive structural improvement arises from the refined behavior of the combined null expansion $\mbox{tr}\chi+\mbox{tr}\underline{\chi}$. Owing to the matching construction and the null structure equations (see (29)), a cancellation occurs near the interface sphere $S_{-a,0}$ which yields the sharper estimate
$|\mbox{tr}\chi+\mbox{tr}\underline{\chi}|=O\!\big(a^{-5/2}\big),$
rather than the individually expected $O(a^{-1})$ bounds. This strengthened control is precisely what enforces the required focusing of the generalized mean curvature $H-|\mbox{tr}_{\Sigma}K|$ and breaks the apparent balance between interior metric loss and collar gain.
The cancellation mechanism described above yields the sharpened bound $|k|\lesssim a^{-3/2}$, consistent with the estimates for the remaining Ricci coefficients and stable under the matching construction. As a consequence, the metric transport estimates improve to
(51) $g\Big(1-\frac{C\epsilon}{a^{3/2}}\Big)\;\leq\;(\Phi_{t}^{-1})^{*}g\;\leq\;g\Big(1+\frac{C\epsilon}{a^{3/2}}\Big)$
on the evolved interior image $\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$, for a universal numerical constant $C>0$. In contrast, the geometric gain contributed by the collar region
$\Big(J^{+}(\mathcal{M}^{1/a}_{\mathrm{int}})\cap\mathcal{M}_{-a-1/a+\epsilon}\Big)\setminus\Phi_{t}(\mathcal{M}^{1/a}_{\mathrm{int}})_{-a-1/a+\epsilon}$
remains of order $O(1)$ and therefore satisfies the lower bound
(52) $\text{Rad}_{\mathrm{collar}}\;\geq\;\text{Rad}(\mathcal{M}^{1/a}_{\mathrm{int}})\Big(1+\frac{C\epsilon}{a}\Big).$
Since $a^{-3/2}\ll a^{-1}$ for $a\gg 1$, the collar contribution dominates the possible metric contraction in the interior. It follows that for sufficiently large $a$ one obtains a strict radius increase,
(53) $\text{Rad}\!\Big(J^{+}(\mathcal{M}_{1})\cap\mathcal{M}_{-a-1/a+\epsilon}\Big)\;\geq\;\text{Rad}(\mathcal{M}_{1})\,\sqrt{1+\frac{1}{10a}},$
after adjusting constants.
Invoking the Schoen–Yau barrier criterion [44], the existence of a MOTS in $J^{+}(\mathcal{M}_{1})\cap\mathcal{M}_{-a-1/a+\epsilon}$ together with the absence of trapped surfaces in $\mathcal{M}_{1}$ is reduced to verifying the sufficient condition
(54) $\frac{3\pi}{2\,\text{Rad}(\mathcal{M}_{1})\sqrt{1+\frac{1}{a}}}\;$ $<\;-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{\prime}|\,|\hat{\underline{\chi}}|^{2}(u^{\prime},\epsilon)\,du^{\prime}$
$\quad+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{\prime}|^{2}}\int_{u_{\infty}}^{u^{\prime}}|u^{\prime\prime}|^{2}\,|\hat{\underline{\chi}}|^{2}(u^{\prime\prime})\,du^{\prime\prime}\,du^{\prime}\;<\;\frac{3\pi}{2\,\text{Rad}(\mathcal{M}_{1})}.$
In particular, if the initial outgoing shear and the magnitude of the incoming expansion at past null infinity $u=u_{\infty}$ are sufficiently large, then the boundary sphere
$\partial\!\Big(J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{t=-a-1/a+\epsilon}\Big)=S_{-a-1/a,3/4}$
acquires sufficiently large generalized mean curvature $H-|\kappa|$. The Schoen–Yau criterion therefore guarantees the existence of a MOTS in the interior of $J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{t=-a-1/a+\epsilon}$.
Let us define the following entity
(55) $\mathsfbf{H}:=-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},0)}{a}+\frac{9}{10a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}+\frac{9}{10a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}$
which is nothing but the generalized mean curvature $(H-|\mbox{tr}_{\Sigma}K|)(-a-1/a,\epsilon)$ for $\epsilon\in[0,1]$ upto negligible $O(a^{-5/2})$ error terms. Before stating the technical version of the main theorem, let us define the locally uniform Sobolev norms.Fix a smooth cutoff $\chi\in C_{c}^{\infty}(B_{2}(0))$ with $\chi\equiv 1$ on $B_{1}(0)$, and for each $y\in\mathcal{M}^{1/a}_{int}$ define $\chi_{y}(x):=\chi(x-y)$. For an integer $s\geq 4$ set
$\|u(t)\|_{H^{s}_{\mathrm{ul}}(\mathcal{M}^{1/a}_{int}))}:=\sup_{y\in\mathcal{M}^{1/a}_{int}}\ \|\chi_{y}\,u(t)\|_{H^{s}(\mathcal{M}^{1/a}_{int})}$
(and similarly for $L^{2}_{\mathrm{ul}},\,L^{\infty}_{\mathrm{ul}}$).
Theorem 2.2 (Main Theorem: Technical Version).